If 4 apples and 2 oranges equals $1 and 2 apples and 3 orange equals $0.70, how much does each apple and each orange cost? There are no quantity discounts.
Answer by stanbon(75887)

If 4 apples and 2 oranges equals $1 and 2 apples and 3 orange equals $0.70, how much does each apple and each orange cost? There are no quantity discounts.
---------------
Equations:
4a + 2r = 100 cents
2a + 3r = 70 cents
--------------------
To eliminate the "r" terms multiply
thru the 1st equation by 3; multiply
thru the 2nd equation by 2.
-----------------------------------------
12a + 6r = 300
4a + 6r = 140
-------------------------
Subract the 2nd from the 1st Eq and solve for "a":
8a = 160
a = 20 cents (price of one apple)
-----------------------------------------
Substitute for "a" in 2a+3r = 70 and solve for "r":
2*20 + 3r = 70
3r = 30
r = 15 cents (price of one orange)
====================================
Cheers,
stan H.